a chemist needs 6 liters of 20 percent acid solution. he has available 25 percent acid solution and a 10 percent acid solution. how much of each solution should be mixed in order to create thsi 20 percent solution
could u list the steps and formula?

Answer by mananth(16946)

percent ---------------- quantity
25 ---------------- x
10 ---------------- 6-x
20.00% ---------------- 6
...
25x+10(6 -x)=20*6

25x+60-10x=120
25x-10x =120-60
15x=60
/15
x=4 liters 25% solution
2 liters 10% solution



2